OpenStudy (anonymous):

Easy yet cute problem, if \( p^{12}=q^6=r^3=s^2, p \neq 1 \). Find the value of \(\log_p pqrs \). Lets see who can solve this the fastest. Pleas read on for the extra question (posted later): How many values of \(x \) satisfy the equation \( \log (2x)=\frac 14 \log (x-15)^4 \)?

OpenStudy (anonymous):

is it 13

OpenStudy (apoorvk):

23/12

OpenStudy (apoorvk):

13

OpenStudy (anonymous):

q=p^2 r=p^4 s=p^6 2+4+6+1=13

OpenStudy (apoorvk):

I think too much actually :/ and act too hastily. :/ It is 13.

OpenStudy (anonymous):

I like posting sitters for a change ... :P

OpenStudy (asnaseer):

I think it can have several values:\[1+4\pm2\pm6\]

OpenStudy (asnaseer):

since q and s can be negative

OpenStudy (anonymous):

Asnaseer, we do not know if the even number radical function is defined for strictly positive or negative. Positive is assumed.

OpenStudy (asnaseer):

so are you saying, for example, that:\[s=-p^6\]is not a valid solution to this?

OpenStudy (anonymous):

Hmm... Rather hard to say exactly what I mean. It's true that \(s=-p^6\). But, I assumed we were using the positive roots only because that's conventional. Otherwise, as you see, our function is actually a multivalued relation rather than a function and thus the logarithm becomes murkily defined. (I am potentially abusing terminology here.)

OpenStudy (asnaseer):

maybe FFM can clarify?

OpenStudy (anonymous):

My main point is that if we allow ourselves to take both of the roots, the logarithm is no longer a function. This can't be the case, since the logarithm is always a function.

OpenStudy (asnaseer):

BTW: You are most probably correct Limitless. I am not a /true/ mathematician as such - I only follow maths as a hobby.

OpenStudy (anonymous):

I hope you realize that I am still impressed by your solution, critique aside.

OpenStudy (asnaseer):

:) thx

OpenStudy (anonymous):

Asnaseer is technically right, but I am happy with the less pedantic solution for this one.

OpenStudy (asnaseer):

Is there some article on this area that I could read about to help me understand this better?

OpenStudy (anonymous):

asnaseer, precisely what confuses you? This is more of a notion of convention rather than of serious mathematical concepts.

OpenStudy (asnaseer):

your description of log being a function and therefore it can only have one value - I was just wondering if there was any formal definition of this concept somewhere?

OpenStudy (apoorvk):

Well. I seem to think that 'q' and 's' can be negative, but their resultant powers would definitely be positive - so how does the question of the powers being 'negative' arise? s= (-p)^6 ---> alrighty! but the exponent can only be '6' right? of the exponent changes sign, then the whole thingy changes. Or am I totally wrong once again? -_-

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Function_(mathematics)#Formal_definition :) I rather love formal definitions if they are also intuitively beautiful. This is one such case. The reason that functions are considered injective and surjective (i.e. one-to-one, which is what I have been saying) is that this allows us to suppose the existence of their inversion via another function! It's rather interesting stuff. :)

OpenStudy (asnaseer):

good point apoorvk!

OpenStudy (asnaseer):

does that mean that:\[\log_p(-p)^6\]is undefined?

OpenStudy (asnaseer):

thx for the link Limitless.

OpenStudy (asnaseer):

sorry I meant:\[\log_p-(p)^6\]

OpenStudy (anonymous):

In real yes.

OpenStudy (anonymous):

hmm.. Good question. \(\log_{p}(-p^6)=\log_{p}(-1)+\log_{p}(p^6)=\log_{p}(-1)+6\) Here's an intuitive way of thinking: What in the world is \(\log_{p}(-1)\)???

OpenStudy (anonymous):

I mean it is not defined in \( \mathbb{R}\)

OpenStudy (asnaseer):

It always amazing me how even /seemingly/ simple problems can generate a whole raft of new learning! :)

OpenStudy (asnaseer):

*amazes

OpenStudy (apoorvk):

Yup^!

OpenStudy (anonymous):

There is a tricky problem on this topic that I encountered last month,I will post it some time later.

OpenStudy (asnaseer):

BTW: The wolf says \(\log(-1)=i\pi\)

OpenStudy (asnaseer):

so FFM - you are right - it does not belong in the reals

OpenStudy (anonymous):

You will find it very intriguing if you ask yourself why we restrict ourselves to certain domains on popular function. It taught me many deep ideas to ponder this. A rather elementary reason is that much of a function cannot even be graphed on a real 2-dimensional plane. Consider \(f(x)=x^2\) at \(x=i\). It is clear that we can graph \(f(i)=-1\), however how do we plot \(i\)??? This is where it becomes very intriguing: Complex numbers require special plotting techniques or special geometric spaces. :)

OpenStudy (asnaseer):

You just melted by brain Limitless! (in a good way) :D

OpenStudy (anonymous):

Well, I am glad I could! :) May someone more enlightened than me correct me if I am misrepresenting things.

OpenStudy (anonymous):

How many values of \(x\) satisfy the equation \( \log (2x)=\frac 14 (x-15)^4 \)?

OpenStudy (lgbasallote):

42 =_=

OpenStudy (asnaseer):

Limitless: That little plotting problem you posed up there is indeed very intriguing. Looks like my weekend is going to be filled with reading material :)

OpenStudy (anonymous):

Good lord, FFM. Surely this does not require numerical methods. D: I hope so, Asn. I had much spare time in class, LOL. You ask yourself questions like this out of boredom.

OpenStudy (anonymous):

Well for complex numbers we need a completely new reference axes, The argand plane.

OpenStudy (anonymous):

*plotting

OpenStudy (asnaseer):

FFM: Can this be solved analytically?

OpenStudy (anonymous):

Oh yes, 1 line if you start sufficiently from left ;)

OpenStudy (asnaseer):

ah! - two values?

OpenStudy (anonymous):

This whole factor \((x-15)^5\) is turning me off.

OpenStudy (anonymous):

No. Btw consider \(\mathbb{R} \)

OpenStudy (asnaseer):

I was thinking in \(\log(2x)\), 2x cannot be negative, therefore x must be positive. And I /assumed/ that \(\frac 14 (x-15)^4\) would have 2 negative and 2 positive solutions, so only 2 solutions possible.

OpenStudy (anonymous):

\(x\) is not an integer. Is this true?

OpenStudy (anonymous):

No.

OpenStudy (experimentx):

\( \log(2x)^4 = (x-15)^4\)

OpenStudy (anonymous):

I missed something, then.

OpenStudy (experimentx):